# CodeEditorBench: Evaluating Code Editing Capability of Large Language Models

### v1

@@ Abstract
These authors contributed equally. Corresponding Authors.
@@ 1 Introduction
  • Highlight the models’ limitations in code polishing and code rewriting as required. While GPT4 shows commendable performance in three out of four areas, its capabilities in code polishing are notably inadequate. Similarly, Gemini Ultra falls short in tasks involving code requirement alterations.
@@ 2 Related Work
Code LLMs  The field witnesses significant growth in developing code LLMs to address the challenges in code understanding and generation. This trend starts with the introduction of Codex (Chen et al., 2021) by OpenAI, followed by the emergence of many influential models including CodeGen (Nijkamp et al., 2023b), CodeT5 (Wang et al., 2021, 2023), and InCoder (Fried et al., 2023). Recent popular open-source models like CodeLLaMa (Rozière et al., 2024), DeepSeek Coder (Guo et al., 2024), and StarCoder (Li et al., 2023a; Lozhkov et al., 2024) represent the forefront of this field. They demonstrate excellent abilities in various code understanding and generation tasks by extensively pre-training from scratch on massive code datasets. Additionally, these base models undergo another phase of instruction tuning, empowering them with better instruction-following capability, leading to significant performance improvements in solving various code-related tasks.
Code Benchmark  Many benchmarks are proposed to compare and evaluate code LLMs. However, these primarily focus on natural language and code generation. HumanEval (Chen et al., 2021) is one of the pioneering and most widely used benchmarks for LLM-based code synthesis, consisting of 164 pairs of Python function signature with docstring and the associated test-cases for correctness checking. Another Python-focused dataset, MBPP (Austin et al., 2021a), is created by crowd-sourcing participants to write in summation 974 programming problems, each of which is comprised of the problem statement (i.e., docstring), the function signature, as well as three test-cases. Beyond Python, there are other benchmarks targeting additional languages such as Spider (SQL) (Yu et al., 2018), HumanEval-X (Zheng et al., 2023) (C++, Javascript and Go), CodeContests (Li et al., 2022) (C++ and Java) and MultiPL-E (Cassano et al., 2022) (extending HumanEval and MBPP to 18 programming languages). Competitive programming benchmarks include LeetCode-Hard Gym (Olausson et al., 2023), which evaluates code generation in multiple languages using LeetCode’s server and the OpenAI gym framework. DebugBench (Tian et al., 2024) advances LLM evaluation by focusing on error correction across diverse programming challenges, from syntax to logical mistakes. EditEval (Hu et al., 2023) assesses LLMs’ ability to understand and execute code editing instructions, measuring how accurately models can modify code based on human-written instructions.
@@ 3.1 Problem Definition
[…]
@@ 4.2 Evaluated Models
We select 19 of the most popular current code LLMs from existing leaderboards evalplus (Liu et al., 2023), bigcode (Ben Allal et al., 2022), encompassing both open-source and closed-source models, with sizes ranging from 6.7B to 34B, including base models and instruction-tuning models. The open-source models include the GPT series (GPT-3.5-Turbo, GPT-4), the Gemini series (Gemini-Pro, Gemini-Ultra), and GLM-4. Closed-source models comprise the CodeLlama series (CL-34B-Base, CL-{7, 13, 34}B-INST), the DeepSeek series (DS-33B-INST), and the outstanding instruction-finetuned models, the WizardCoder series (WC-15B based on StarCoder, WC-33B based on DS-33B-Base), OctoCoder based on StarCoder, CF-34B based on CL-34B-Python, PH-34B based on CL-34B-Base, the Magicoder series (MC-DS-6.7B based on DS-6.7B-Base, MC-CL-7B based on CL-7B) and the OpenCodeInterpreter series(OpenCI-DS-6.7B based on DS-6.7B-Base, OpenCI-DS-33B based on DS-33B-Base). Appendix B presents more detailed information regarding our evaluated models.
@@ Appendix B Evaluated Models
ise-uiuc/Magicoder-S-DS-6.7B 6.7B 2023-12-04 Magicoder-S-DS-6.7B
ise-uiuc/Magicoder-S-CL-7B 7B 2023-12-04 Magicoder-S-CL-7B
bigcode/octocoder 15.5B 2023-08-14 octocoder
WizardLM/WizardCoder-15B-V1.0 15B 2023-06-16 WizardCoder-15B-V1.0
WizardLM/WizardCoder-33B-V1.1 33B 2024-01-04 WizardCoder-33B-V1.1
deepseek-ai/deepseek-coder-33b-instruct 33B 2023-02-01 deepseek-coder-33b-instruct
codefuse-ai/CodeFuse-CodeLlama-34B 34B 2023-09-11 CodeFuse-CodeLlama-34B
Phind/Phind-CodeLlama-34B-v2 34B 2023-09-01 Phind-CodeLlama-34B-v2
m-a-p/OpenCodeInterpreter-DS-6.7B 6.7B 2024-02-22 OpenCodeInterpreter-DS-6.7B
m-a-p/OpenCodeInterpreter-DS-33B 33B 2024-02-22 OpenCodeInterpreter-DS-33B
codellama/CodeLlama-34b-hf 34B 2023-08-24 CodeLlama-34b-hf
codellama/CodeLlama-7b-Instruct-hf 7B 2023-08-24 CodeLlama-7b-Instruct-hf
codellama/CodeLlama-13b-Instruct-hf 13B 2023-08-24 CodeLlama-13b-Instruct-hf
codellama/CodeLlama-34b-Instruct-hf 34B 2023-08-24 CodeLlama-34b-Instruct-hf
gpt-3.5-turbo-1106 - 2021-10-01 gpt-3.5-turbo-1106
gpt-4-0613 - 2021-10-01 gpt-4-0613
glm-4 - 2024-01-16 glm-4
gemini-pro - 2023-12-06 gemini-pro
gemini-ultra - 2023-12-06 -
@@ Appendix E Pass Criteria
Debug 300s for all test cases. 30s for single test case. 512512512MB Pass all test cases.
@@
Polish $\bar{T}$ms $\bar{M}$ Pass all test cases and $\bar{T}_{\text{avg}}<\bar{T}$ or $\bar{M}_{\text{avg}}<\bar{M}$

### v2

@@ Abstract
${}^{\ast}$ These authors contributed ${}^{{\dagger}}$ Corresponding Authors.
@@ 1 Introduction
  • Highlight the models’ limitations in code polishing and code rewriting as required. While GPT-4 shows commendable performance in three out of four areas, its capabilities in code polishing are notably inadequate. Similarly, Gemini Ultra falls short in tasks involving code requirement alterations.
@@ 2 Related Work
Code LLMs  The field witnesses significant growth in developing code LLMs to address the challenges in code understanding and generation. This trend starts with the introduction of Codex (Chen et al., 2021) by OpenAI, followed by the emergence of many influential models including CodeGen (Nijkamp et al., 2023b), CodeT5 (Wang et al., 2021, 2023), and InCoder (Fried et al., 2023). Recent popular open-source models like CodeLLaMa (Rozière et al., 2024), DeepSeek Coder (Guo et al., 2024), and StarCoder (Li et al., 2023a; Lozhkov et al., 2024) represent the forefront of this field. They demonstrate excellent abilities in various code understanding and generation tasks by extensively pre-training from scratch on massive code datasets. Additionally, these base models undergo another phase of instruction tuning (Zheng et al., 2024; Wei et al., 2023; Luo et al., 2023; Royzen et al., 2023; Liu et al., 2023a; Muennighoff et al., 2024), empowering them with better instruction-following capability, leading to significant performance improvements in solving various code-related tasks.
Code Benchmark  Many benchmarks are proposed to compare and evaluate code LLMs. However, these primarily focus on natural language and code generation. HumanEval (Chen et al., 2021) is one of the pioneering and most widely used benchmarks for LLM-based code synthesis, consisting of 164 pairs of Python function signature with docstring and the associated test-cases for correctness checking. Another Python-focused dataset, MBPP (Austin et al., 2021a), is created by crowd-sourcing participants to write in summation 974 programming problems, each of which is comprised of the problem statement (i.e., docstring), the function signature, as well as three test-cases. Beyond Python, there are other benchmarks targeting additional languages such as Spider (SQL) (Yu et al., 2018), HumanEval-X (Zheng et al., 2023) (C++, Javascript and Go), HumanEvalPack (Muennighoff et al., 2024)(Python, JavaScript, Java, Go, C++ and Rust), CodeContests (Li et al., 2022) (C++ and Java) and MultiPL-E (Cassano et al., 2022) (extending HumanEval and MBPP to 18 programming languages). Competitive programming benchmarks include LeetCode-Hard Gym (Olausson et al., 2023), which evaluates code generation in multiple languages using LeetCode’s server and the OpenAI gym framework. DebugBench (Tian et al., 2024) advances LLM evaluation by focusing on error correction across diverse programming challenges, from syntax to logical mistakes. EditEval (Hu et al., 2023) assesses LLMs’ ability to understand and execute code editing instructions, measuring how accurately models can modify code based on human-written instructions.
@@ 3.1 Problem Definition
  • an ideal debugger D that rectifies any buggy code from $c$ to $*c*$ should satisfy that D($c$) = $*c*$ s.t.$\forall$($x_{i}$, $y_{i}$) $\in R_{i}$, $*a_{c}*$($x_{i}$) = $y_{i}$. Debugging can be regarded as the converting process of debugger D.
  • an ideal translater T that translates any code from language a to language b should satisfy that T($c$) = $*c*$ s.t. $\forall$($x_{i}$,$y_{i}$) $\in R_{i}$, $a_{c}$($x_{i}$) = $y_{i}$ and $*a_{c}*$($x_{i}$) = $y_{i}$. Translating can be regarded as the converting process of translater T.
  • an ideal optimizer P that make any code more effective in either time complexity or space complexity with the guarantee of accuracy should satisfy that P($c$) = $*c*$ s.t. $\forall$($x_{i}$,$y_{i}$) $\in R_{i}$, $a_{c}$($x_{i}$) = $y_{i}$,$*a_{c}*$($x_{i}$) = $y_{i}$ and $avg_{time}(c*)\leq avg_{time}(c)$ or $avg_{memory}(c*)\leq avg_{memory}(c)$. Polishing can be regarded as the converting process of optimizer P
  • an ideal requirement switcher S that switches a similar code that satisfies question $a$ to a target code that satisfies question $b$ based on the sample inputs and outputs of $a$ and $b$ should satisfy that S($c$) = $*c*$ s.t. $\forall$($x_{i}$,$y_{i}$) $\in R_{b}$, $*a_{c}*$($x_{i}$) = $y_{i}$.Switching can be regarded as the converting process of switcher S.
@@ 4.2 Evaluated Models
We select 19 of the most popular current code LLMs from existing leaderboards evalplus (Liu et al., 2023b), bigcode (Ben Allal et al., 2022), encompassing both open-source and closed-source models, with sizes ranging from 6.7B to 34B, including base models and instruction-tuning models. The open-source models include the GPT series (GPT-3.5-Turbo, GPT-4), the Gemini series (Gemini-Pro, Gemini-Ultra), and GLM-4. Closed-source models comprise the CodeLlama series (CL-34B-Base, CL-{7, 13, 34}B-INST), the DeepSeek series (DS-33B-INST), and the outstanding instruction-finetuned models, the WizardCoder series (WC-15B based on StarCoder, WC-33B based on DS-33B-Base), OctoCoder based on StarCoder, CF-34B based on CL-34B-Python, PH-34B based on CL-34B-Base, the Magicoder series (MC-DS-6.7B based on DS-6.7B-Base, MC-CL-7B based on CL-7B) and the OpenCodeInterpreter series(OpenCI-DS-6.7B based on DS-6.7B-Base, OpenCI-DS-33B based on DS-33B-Base). Appendix B presents more detailed information regarding our evaluated models.
@@ Appendix B Evaluated Models
ise-uiuc/Magicoder-S-DS-6.7B
6.7B 2023-12-04
Magicoder-S-DS-6.7B
ise-uiuc/Magicoder-S-CL-7B
7B 2023-12-04
Magicoder-S-CL-7B
bigcode/octocoder
15.5B 2023-08-14
octocoder
WizardLM/WizardCoder-15B-V1.0
15B 2023-06-16
WizardCoder-15B-V1.0
WizardLM/WizardCoder-33B-V1.1
33B 2024-01-04
WizardCoder-33B-V1.1
deepseek-ai/deepseek-coder-33b-instruct
33B 2023-02-01
deepseek-coder-33b-instruct
codefuse-ai/CodeFuse-CodeLlama-34B
34B 2023-09-11
CodeFuse-CodeLlama-34B
Phind/Phind-CodeLlama-34B-v2
34B 2023-09-01
Phind-CodeLlama-34B-v2
m-a-p/OpenCodeInterpreter-DS-6.7B
6.7B 2024-02-22
OpenCodeInterpreter-DS-6.7B
m-a-p/OpenCodeInterpreter-DS-33B
33B 2024-02-22
OpenCodeInterpreter-DS-33B
codellama/CodeLlama-34b-hf
34B 2023-08-24
CodeLlama-34b-hf
codellama/CodeLlama-7b-Instruct-hf
7B 2023-08-24
CodeLlama-7b-Instruct-hf
codellama/CodeLlama-13b-Instruct-hf
13B 2023-08-24
CodeLlama-13b-Instruct-hf
codellama/CodeLlama-34b-Instruct-hf
34B 2023-08-24
CodeLlama-34b-Instruct-hf
gpt-3.5-turbo-1106 (OpenAI, 2022)
- 2021-10-01
gpt-3.5-turbo-1106
gpt-4-0613 (OpenAI, 2023)
- 2021-10-01
gpt-4-0613
glm-4 (zhipuai, 2024)
- 2024-01-16
glm-4
gemini-pro (Google, 2023)
- 2023-12-06
gemini-pro
gemini-ultra (Google, 2023)
- 2023-12-06
@@ Appendix E Pass Criteria
Debug 300s for all test cases. 30s for single test case. $512$MB Pass all test cases.
@@
Polish $\bar{T}$ms $\bar{M}$
Pass all test cases
and
$\bar{T}_{\text{avg}}<\bar{T}$ or $\bar{M}_{\text{avg}}<\bar{M}$